# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

### v1

@@ 3 Bilevel model and Zeroth-order-first-order method
1:  Input: $\eta>0$, $\zeta>0$, ${\mathbf{s}}^{0}={\mathbf{s}}^{k}$, $K,T\in\mathbb{N}_{+}$, $\lambda\geq 0$.
2:  for k=0,…,K do
3:     for $t=0,\dots,T-1$do
4:        Let ${\mathbf{s}}^{k}_{t+1}={\mathbf{s}}^{k}_{t}-\eta\nabla_{\mathbf{{\mathbf{s}}}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k}_{t})$.
5:        Output ${\mathbf{s}}^{k+1}={\mathbf{s}}^{k}_{T}$.
@@
7:     Let ${\bm{\theta}}^{k+1}={\bm{\theta}}^{k}-\zeta\nabla_{\bm{\theta}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k})$ and ${\mathbf{p}}^{k+1}={\mathbf{p}}^{k}-\zeta\nabla_{\mathbf{p}}G_{\lambda}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k}).$
@@
However, note that
$\nabla_{\bm{\theta}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k})=\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})+\\ \lambda_{k}(\nabla_{{\bm{\theta}}}F({\bm{\theta}},{\mathbf{p}}^{k};\mathfrak{D}_{\mathbf{p}})+\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k};\mathfrak{D}_{\mathbf{p}})),$ (6)
requires calculating the gradient with respect to ${\bm{\theta}}$, i.e, $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})$. Given the large scale of ${\bm{\theta}}$ in LLMs, this is computationally expensive. To avoid this, we use zeroth-order (ZO) information to approximate the gradient $\nabla_{\bm{\theta}}G$. Following (Malladi et al., 2023; Zhang et al., 2024b; Guo et al., 2024b), we employ the Simultaneous Perturbation Stochastic Approximation (SPSA) as a classical zeroth-order gradient estimator. Specifically, at each iteration $k$, we sample ${\mathbf{z}}^{k}\sim N(0,I_{d})$, where $d$ is the dimension of ${\bm{\theta}}$. We then approximate the gradient $\nabla_{\bm{\theta}}F$ as follows:
^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x):=\\ \frac{F({\bm{\theta}}^{k}+\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)-F({\bm{\theta}}^{k}-\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)}{2\epsilon}{\mathbf{z}}^{k}.$ (7)
Additionally, in each iteration $k$ and inner iteration $t$ of Algorithm 1, we sample a mini-batch $\mathcal{B}$ of size $B$. We use ^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B})$ to substitude $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{f})$ and $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{p})$in (6). We also use mini-batches when calculating the gradients with respect to the PEFT parameters ${\mathbf{s}}$ and ${\mathbf{p}}$.
This approach leads to the final algorithm (Algorithm 2 and Figure 1) for fine-tuning LLMs using the bilevel model (2). We refer to this method as the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method.
1:  Input: $\eta>0$, $\zeta>0$, batchsize $B$, ${\mathbf{s}}^{0}={\mathbf{s}}^{k}$, $K,T\in\mathbb{N}_{+}$, $\lambda>0$.
2:  for k=0,…,K do
3:     for t=0,…,T-1 do
4:        Sample a batch $\mathcal{B}^{k}_{t,{\mathbf{p}}}$ from $\mathfrak{D}_{\mathbf{p}}$.
5:        Let ${\mathbf{s}}^{k}_{t+1}={\mathbf{s}}^{k}_{t}-\eta\nabla_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k}_{t};\mathcal{B}^{k}_{t,{\mathbf{p}}})$
6:        Output ${\mathbf{s}}^{k+1}={\mathbf{s}}^{k}_{T}$.
7:     end for
8:     Sample a batch $\{\mathcal{B}^{k}_{f}\}$ from $\mathfrak{D}_{f}$ and $\{\mathcal{B}^{k}_{\mathbf{p}}\}$ from $\mathfrak{D}_{\mathbf{p}}$.
9:     For $x\in\mathcal{B}^{k}_{\mathbf{p}}\cup\mathcal{B}^{k}_{f}$, calculate ^$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x)$ following (7).
10:     Let
${\mathbf{p}}^{k+1}={\mathbf{p}}^{k}-\zeta(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\\ \lambda_{k}(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})))$ (8)
${\bm{\theta}}^{k+1}={\bm{\theta}}^{k}-\zeta(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\\ \lambda_{k}(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})-\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k+1};\mathcal{B}^{k}_{\mathbf{p}})))$ (9)
11:  end for
@@ 4 Experiments
We conduct extensive experiments on various LLMs of different scales to demonstrate the effectiveness of bilevel-ZOFO in both single-task and multi-task meta-training settings. Our proposed structure is able to incorporate any variation of zeroth-order methods in the upper-level step and any PEFT method in the lower level. To maintain focus on testing the effectiveness of the proposed bilevel structure and its unique multitask learning capabilities, we used the classic MeZO  (Malladi et al., 2023).
4.1 Bilevel-ZOFO for Single-task Fine-tuning
Following MeZO (Malladi et al., 2023), we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task. We employ prompt-tuning (Lester et al., 2021), prefix-tuning (Li & Liang, 2021), and LoRA (Hu et al., 2022)- well-known PEFT baselines-for lower-level training to validate bilevel-ZOFO under different conditions and resource constraints. During each lower-level update, we update only the PEFT parameters, and during the upper-level optimization step, we tune the full model using zeroth-order gradient approximation. We perform 10 lower-level updates between each pair of upper-level updates. For each task, we randomly sample 1000 examples for training, 500 examples for validation, and 1000 examples for testing. We use the Adam optimizer (Kingma & Ba, 2015) and report test accuracy or F1-score.
We compare our method against several baselines, including MeZO for Full Model Fine-tuning, MeZO for PEFT, and First-order PEFT. We fix the total memory budget of each step across bilevel-ZOFO and the baselines. We train zeroth-order methods for 10,000 steps, and first-order methods for 5000 steps. For all experimental details, refer to the Appendix B.1.3 and Appendix B.1.4. We also provide the training loss for the lower-level objective of the bilevel framework in Figure 4 of Appendix B.2 to show that consistent with the guarantees provided by our theoretical analysis in Section 5, Bilevel-ZOFO converges. We compare the memory requirements of our method with the baselines in Figure 5, and provide wall-clock analysis in Table 7. Figure 5 and Table 7 show that bilevel-ZOFO does not increase the memory usage of PEFT and MeZO.
Bileve-ZOFO mitigates sensitivity to hard prompts: We begin our experiments by showing that Bilevel-ZOFO mitigates the sensitivity of MeZO to hard prompts. First, similar to Table 5 of the MeZO (Malladi et al., 2023) paper, we experiment with different choices of hard prompts for both MeZO and our Bilevel-ZOFO. Table 1 shows the results for tuning Opt 1.3b on SST2 and COPA. We can see that our method effectively mitigates the sensitivity of MeZO to hard prompts. The difference between the results with and without a simple hard prompt in our experiment is much less than MeZO’s. See Appendix B.4 for a full discussion.
@@ 4.1 Bilevel-ZOFO for Single-task Fine-tuning
Bilevel-ZOFO SST-2 93.3 92.9 -0.4
COPA 76.66 73.6 -3.06
Table 1: Prompt Sensitivity Comparison for MeZO and Bilevel-ZOFO. Bilevel-ZOFO effectively mitigates the sensitivity of MeZO to hard prompts.
@@
Table 2 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B (Zhang et al., 2022) on a downstream task. Table 3 demonstrates the results for Llama2-7b (Touvron et al., 2023). We can make the following observations:
Bilevel-ZOFO outperforms MeZO on almost all tasks: With the same memory allocation per training step, bilevel-ZOFO outperforms MeZO, even when trained for half the number of iterations across most tasks.
Bilevel-ZOFO outperforms FO PEFT on average:
Comparing each FO-PEFT setting with the corresponding bilevel-ZOFO setting, we see that bilevel-ZOFO outperforms the corresponding FO-PEFT methods across most instances and on average.
Bilevel-ZOFO outperforms baselines more significantly in resource-constrained settings: The number of parameters tuned for prefix tuning and prompt tuning is lower than for LoRA (See Figure 3). As shown in Table 2, when fewer parameters are tuned, bilevel-ZOFO demonstrates a larger improvement over first-order methods in tuning PEFT models. Since memory usage and training steps remain the same, bilevel-ZOFO proves to be a more suitable option for fine-tuning LLMs in constrained environments compared to PEFT and MeZO.
Bilevel-ZOFO scales effectively to larger LLMs: Table 3 compares bilevel-ZOFO with the baselines when fine-tuning Llama2-7b (Touvron et al., 2023) on various classification and open-ended generation tasks. The results show that bilevel-ZOFO’s advantages are not confined to smaller models like OPT-1.3b, but also extend to larger LLMs.
4.2 Multi-Task Fine-Tuning Experiments
Following Min et al. (2022), we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. For instance, in non-classification-to-classification setting, we train on a number of non-classification subtasks and test on a number of distinct classification subtasks. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT (Ye et al., 2021) and UNIFIEDQA (Khashabi et al., 2020), comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification and question answering all in English. We use GPT2-Large (Radford et al., 2019) as the base model for these experiments.
We compare our method against several baseline approaches:
  • MetaICL (Min et al., 2022): A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, ${(a_{1},b_{1}),\dots,(a_{k},b_{k})}$, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
We sample 768 examples from each training sub-task. We train MetaICL in their original setting for 30,000 steps. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we DO NOT include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  (Malladi et al., 2023), we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on 4 demonstration samples from each class of each test sub-task. Similar to (Min et al., 2022), we report Macro-averaged F1 as the evaluation metric. See Appendix B.5 for all training details.

### v2

@@ 3.2 Bilevel Model
1:Symbols: $\bm{\theta}\in\mathbb{R}^{d}$ (backbone params), $\mathbf{p}\in\mathbb{R}^{d^{\prime}}$ (PEFT params), $\mathbf{s}\in\mathbb{R}^{d^{\prime}}$ (aux inner variable), $G_{\lambda}(\bm{\theta},\mathbf{p},\mathbf{s})$ (penalized objective), $K$ (number of outer steps), $T$ (number of inner steps between two outer steps), $\eta>0$ (inner LR), $\zeta>0$ (outer LR), $\{\lambda_{k}\}\subseteq\mathbb{R}_{+}$ (penalty at step $k$).
2:Input: step sizes $\eta,\zeta>0$; initial states $\bm{\theta}^{0},\mathbf{p}^{0},\mathbf{s}^{0}$; $K,T\in\mathbb{N}_{+}$; penalty schedule $\{\lambda_{k}\}_{k=0}^{K-1}$.
3:for $k=0,\dots,K-1$ do
4:  for $t=0,\dots,T-1$ do
5:   $\mathbf{s}^{k}_{t+1}=\mathbf{s}^{k}_{t}-\eta\,\nabla_{\mathbf{s}}\,G_{\lambda_{k}}(\bm{\theta}^{k},\mathbf{p}^{k},\mathbf{s}^{k}_{t})$
@@
7:  $\mathbf{s}^{k+1}\leftarrow\mathbf{s}^{k}_{T}$
8:  $\bm{\theta}^{k+1}=\bm{\theta}^{k}-\zeta\,\nabla_{\bm{\theta}}\,G_{\lambda_{k}}(\bm{\theta}^{k},\mathbf{p}^{k},\mathbf{s}^{k+1})$
9:  $\mathbf{p}^{k+1}=\mathbf{p}^{k}-\zeta\,\nabla_{\mathbf{p}}\,G_{\lambda_{k}}(\bm{\theta}^{k},\mathbf{p}^{k},\mathbf{s}^{k+1})$
10:end for
In this way, we can solve the bilevel problem as a minimax problem. The basic minimax algorithm works as follows: at iteration $k$, we first solve the maximization problem $\max_{\mathbf{{\mathbf{s}}}}G_{\lambda}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}})$ with $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ fixed. For example, we can update ${\mathbf{s}}^{k}$ using an inner loop with stochastic gradient descent (SGD). Let ${\mathbf{s}}^{k+1}$ be the result of this inner loop. Then, in the outer loop, we update $({\bm{\theta}}^{k},{\mathbf{p}}^{k})$ by solving $\min_{{\bm{\theta}},{\mathbf{p}}}G_{\lambda}({\bm{\theta}},{\mathbf{p}},{\mathbf{s}}^{k+1})$ with ${\mathbf{s}}^{k+1}$ fixed. Again, SGD can be used to update ${\bm{\theta}}^{k}$ and ${\mathbf{p}}^{k}$. The conceptual algorithm is presented in Algorithm 1. We assume we do a total of $K$ outer iterations and $T$ inner iterations between each two consecutive outer steps.
However, note that
$\nabla_{\bm{\theta}}G_{\lambda_{k}}({\bm{\theta}}^{k},{\mathbf{p}}^{k},{\mathbf{s}}^{k})=\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})+\lambda_{k}(\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{\mathbf{p}})+\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k};\mathfrak{D}_{\mathbf{p}})),$ (6)
requires calculating the gradient with respect to $\bm{\theta}$}, i.e, $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathfrak{D}_{f})$. Given the large scale of $\bm{\theta}$} in LLMs, this is computationally expensive. To avoid this, we use zeroth-order (ZO) information to approximate the gradient $\nabla_{\bm{\theta}}G$. Following [33, 61, 14], we employ the Simultaneous Perturbation Stochastic Approximation (SPSA) as a classical zeroth-order gradient estimator. Specifically, at each iteration $k$, we sample ${\mathbf{z}}^{k}\sim N(0,I_{d})$, recalling that $d$ is the dimension of $\bm{\theta}$}. We then approximate the gradient $\nabla_{\bm{\theta}}F$ as follows:
$\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x):=\frac{F({\bm{\theta}}^{k}+\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)-F({\bm{\theta}}^{k}-\epsilon{\mathbf{z}}^{k},{\mathbf{p}}^{k};x)}{2\epsilon}{\mathbf{z}}^{k}.$ (7)
Additionally, in each iteration $k$, we sample mini-batches $\mathcal{B}_{f}^{k}$ and $\mathcal{B}_{p}^{k}$ and use $\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B})$ to substitude $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{f})$ and $\nabla_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{D}_{p})$in (6). We also use mini-batches when calculating the gradients with respect to the PEFT parameters $\mathbf{s}$} and $\mathbf{p}$}.
This approach leads to the final algorithm (Algorithm 2 and Figure 5) for fine-tuning LLMs using the bilevel model (2). We refer to this method as the Bilevel Zeroth-Order-First-Order (Bilevel ZOFO) method. In Appendix B, we show that Bilevel ZOFO converges at a rate of $O(\epsilon^{-2})$ under widely accepted assumptions. [^2] In our theorem, we assume strong convexity on the lower level objective function, as is common in other theoretical work on bilevel optimization. While strong convexity facilitates theoretical analysis, our experiments demonstrate that the method remains robust even when this condition is not strictly satisfied. The complexity of Bilevel ZOFO matches that in previous ZO minimax algorithm in [53] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
1:Symbols: $\bm{\theta}\!\in\!\mathbb{R}^{d}$ (backbone params), $\mathbf{p}\!\in\!\mathbb{R}^{d^{\prime}}$ (PEFT params), $\mathbf{s}\!\in\!\mathbb{R}^{d^{\prime}}$ (aux inner variable), $F(\bm{\theta},\mathbf{p};\mathcal{D})$ (avg. loss over data $\mathcal{D}$), $\widehat{\nabla}_{\bm{\theta}}F$ (ZO grad. estimator; see (7)), $\mathfrak{D}_{\mathbf{p}}$ (inner dataset), $\mathfrak{D}_{f}$ (outer dataset), $B$ (mini-batch size), $K$ (number of outer steps), $T$ (number of inner steps between two outer steps), $\eta>0$ (inner LR), $\zeta>0$ (outer LR), $\{\lambda_{k}\}\subseteq\mathbb{R}_{+}$ (penalty schedule).
2:Input: step sizes $\eta,\zeta>0$; batch size $B$; datasets $\mathfrak{D}_{\mathbf{p}},\mathfrak{D}_{f}$; initial states $\bm{\theta}^{0},\mathbf{p}^{0},\mathbf{s}^{0}$; $K,T\in\mathbb{N}_{+}$; penalty $\{\lambda_{k}\}_{k=0}^{K-1}$.
3:for k=0,…,K do
4:  for t=0,…,T-1 do
5:   Sample a batch $\mathcal{B}^{k}_{t,{\mathbf{p}}}$ from $\mathfrak{D}_{\mathbf{p}}$.
6:   Let ${\mathbf{s}}^{k}_{t+1}={\mathbf{s}}^{k}_{t}-\eta\nabla_{\mathbf{{\mathbf{s}}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k}_{t};\mathcal{B}^{k}_{t,{\mathbf{p}}})$
7:   Output ${\mathbf{s}}^{k+1}={\mathbf{s}}^{k}_{T}$.
@@
9:  Sample a batch $\{\mathcal{B}^{k}_{f}\}$ from $\mathfrak{D}_{f}$ and $\{\mathcal{B}^{k}_{\mathbf{p}}\}$ from $\mathfrak{D}_{\mathbf{p}}$.
10:  For $x\in\mathcal{B}^{k}_{\mathbf{p}}\cup\mathcal{B}^{k}_{f}$, calculate $\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};x)$ following (7).
11:  Let
${\mathbf{p}}^{k+1}={\mathbf{p}}^{k}-\zeta(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\lambda_{k}(\nabla_{{\mathbf{p}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})))$ (8)
${\bm{\theta}}^{k+1}={\bm{\theta}}^{k}-\zeta(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{f})+\lambda_{k}(\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{p}}^{k};\mathcal{B}^{k}_{\mathbf{p}})-\hat{\nabla}_{{\bm{\theta}}}F({\bm{\theta}}^{k},{\mathbf{s}}^{k+1};\mathcal{B}^{k}_{\mathbf{p}})))$ (9)
12:end for
@@ 4 Experiments
We conduct extensive experiments on various LLMs of different scales to demonstrate the effectiveness of bilevel-ZOFO in improving current zeroth order methods and PEFT. We also conduct experiments in testing its potential in meta training. Noticing that our proposed structure is able to incorporate any variation of zeroth-order methods in the upper-level step and any PEFT method in the lower level, to maintain focus on testing the effectiveness of the proposed bilevel structure and its unique multitask learning capabilities, we used the classic MeZO  [33].
4.1 Single Task Experiments
4.1.1 Experimental Setting
Following MeZO [33], we evaluate our approach on a range of classification and multiple-choice tasks. In this setting, training and testing are conducted on the same task. We employ prompt-tuning [22], prefix-tuning [24], and LoRA [16]- well-known PEFT baselines-for lower-level training to validate bilevel-ZOFO under different conditions and resource constraints. During each lower-level update, we update only the PEFT parameters, and during the upper-level optimization step, we tune the full model using zeroth-order gradient approximation. We perform 10 lower-level updates between each pair of upper-level updates. For each task, we randomly sample 1000 examples for training, 500 examples for validation, and 1000 examples for testing. We use the Adam optimizer [20] and report test accuracy or F1-score.
We compare our method against several baselines, including MeZO for Full Model Fine-tuning, MeZO for PEFT, and First-order PEFT. We fix the total memory budget of each step across bilevel-ZOFO and the baselines. We train zeroth-order methods for 10,000 steps, and first-order methods for 5000 steps. For all experimental details, refer to the Appendix D.1. We also provide the training loss for the lower-level objective of the bilevel framework in Figure 2 to show that consistent with the guarantees provided by our theoretical analysis in Section B, Bilevel-ZOFO converges. See Appendix D.2 for more details.
4.1.2 Results
Bilevel-ZOFO mitigates MeZO’s sensitivity to task prompts: We present experimental results demonstrating that Bilevel-ZOFO significantly reduces the prompt sensitivity observed in MeZO.
@@ 4.1.2 Results
Ours SST-2 93.3 92.9 -0.4
COPA 76.7 73.6 -3.1
Table 1: Prompt sensitivity comparison for MeZO and Bilevel-ZOFO. Bilevel-ZOFO effectively mitigates the extreme sensitivity of MeZO to adding task prompts to inputs.
Following the setup of Table 5 in the MeZO paper [33], we evaluate both MeZO and Bilevel-ZOFO in two scenarios: 1- where a simple task prompt is prepended to each input versus 2- where no such prompt is used. Table 1 reports results for tuning OPT-1.3B on SST-2 and COPA using LoRA as the PEFT method. Our findings show that Bilevel-ZOFO is markedly less sensitive to prompt variations than MeZO; the performance gap between prompted and unprompted settings is substantially smaller for Bilevel-ZOFO.
@@
Table 2 presents the test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B [60] on a downstream task. Table 3 demonstrates the results for Llama2-7b [50]. We can make the following observations:
Bilevel-ZOFO offers a better training speed - accuracy tradeoff than MeZO Bilevel-ZOFO outperforms MeZO, even when trained for half the number of iterations across almost all tasks, thus offering a better training duration-performance trade-off than MeZO (Also see Figure 1) .
Bilevel-ZOFO outperforms FO PEFT on most tasks and on average: From Table 2 and Table 3, we see that bilevel-ZOFO outperforms the corresponding FO-PEFT methods across most instances and on average, comparing each FO PEFT setting with the corresponding bilevel-ZOFO setting. This is while using the same level of memory as FO PEFT.
Bilevel-ZOFO scales effectively to larger LLMs: Figure 1 and Table 3 shows that bilevel-ZOFO’s advantages are not confined to smaller models like OPT-1.3b, but also extend to larger LLMs.